We think you have liked this presentation. If you wish to download it, please recommend it to your friends in any social system. Share buttons are a little bit lower. Thank you!
Presentation is loading. Please wait.
Published byBasil Floyd
Modified over 4 years ago
TDC2/TCC2 consolidation EATM 15/07/2014
EN/EL – DC and control cable identification and cut completed – Old light removal in TDC2 – New AUG installed
Demin. circuit valve replacement Remplacement des vannes terminé
Unités de ventilation Travaux terminés.
XTAX upgrade 12/12 XTAX tables XTAX repaired Still alignment issue for XTAX T4 This week: survey & test on XTAX T6 + roof closure
Magnet Manifolds re-installed
OTND (Onet nucleaire) Nettoyage 3 eme chateau XTAX OK Démontage passerelle et chemins de câbles en cours
OTND (Onet nucleaire)
Dose Most critical activitiesCollective dose in uSv 471594186 XTAX table upgrade4186 374052554 Removal, temporary storage and reinstall...2554 197361705 TDC2/TCC2 replace valves1705 438701630 Mesures topographiques TCC21630 447161526 Remplacement flexibles aimants TDC21526 483481104 démantèlement et nettoyage TDC21104 486641063 TDC2 Consolidation works EN/EL/CF - Phas...1063 TOTAL 25000
TCC2/TDC2 consolidation WG ALARA COMMITTEE PREPARATION for Xmas stop activities S. Evrard T. Wijnands
NA YETS: schedule week by week S. Evrard.
Study report on the TDC2/TCC2 Cabling and Low Voltage activities Ivan Moya Martinez– EN/EL/CF – WP 31st Meeting
Lau Gatignon, EATM, 7 October The T9 and T10 lines have been running smoothly from mid July and the T11 beam started up very smoothly for CLOUD.
HiRadMat : report from WG2&3 – WANF dismantling January 15, 2010S. Evrard 1 Worksite: activities achieved or in progress WANF cleaning: status Proposal.
G2P Target installation and Beyond. Schedule Pivot before.
CURRENT STATUS OF CV INSTALLATIONS, WORKS DURING LS1 AND PROBLEMS ENCOUNTERED DURING START-UP NORTH AREA Bill Bannister / Jani Lehtinen EATM – 7 th October.
LIU CABLING & OPTICAL FIBRES STRATEGY Georgi Minchev Georgiev EN-EL-CF.
YETS ACTIVITIES SPS-NA S. Girod & S. Evrard on behalf of the SBA team, EATM,
ELENA Installation follow-up 21/05/2015 Worksite follow-up.
TDC2-TCC2 consolidation status report to EATM EDMS October 7 th, 2014 S. EVRARD, EN-MEF, on behalf of the TDC2/TCC2 WG.
S. Evrard — EN/MEF IEFC Workshop, March 08, 2012 TCC2/TDC2 Shutdown Activities and Consolidation Plans On behalf of the TDC2/TCC2.
08/24/07Shutdown’07 Week 3 F.G. Garcia1 MTA beam line installation Weekly Report August 24, 2007 Period covered: August 17 th to August 24 th (Week 3)
LHC Machine Status Report Roberto Saban LHCC June 4 th 2014.
TDC2-TCC2 consolidation status report to RSSO EDMS October 14 th, 2014 E. Harrouch, Q. Bouirek, EN-MEF, on behalf of the TDC2/TCC2 WG.
Status of handling studies for PSB-LIU Caterina Bertone EN-HE LIU-PSB 17/09/2015 Courtesy: A. Newborough EDMS
- Hearing Session - Marion Picard - July 28 th, 2011 S TATUS OF H I R AD M AT S AFETY F ILE Outline Safety file Descriptive part Hazard inventory Demonstrative.
Injector De-cabling Project Enlarged meeting # E. HARROUCH, EN-MEF, on behalf of the Injector De-cabling Project Team.
© 2019 SlidePlayer.com Inc. All rights reserved.